ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 57 Armas-Sanabria, Lorena, Instituto de Matematicas, UNAM, Mexico 3-Pure Braids and Surgery Let L be the set of links obtained as the closure of 3-pure braids. We are interested in knowing when it is possible to obtain a simply connected 3manifold by integral Dehn surgery (m, n, p) on an element of L, and then in determining if such manifold is S^{3}. Let L be any link in L, we calculate the μ-invariant of the Z_{2}-homology spheres and the Casson’s invariant of the Z-homology spheres obtained by surgery on L. Also, we will see that it is not possible to obtain a simply connected 3-manifold on some families of L.